# REN: Anatomically-Informed Mixture-of-Experts for Interstitial Lung Disease Diagnosis

### v1

@@ Abstract
Mixture-of-Experts (MoE) architectures have significantly contributed to scalable machine learning by enabling specialized subnetworks to tackle complex tasks efficiently. However, traditional MoE systems lack domain-specific constraints essential for medical imaging, where anatomical structure and regional disease heterogeneity strongly influence pathological patterns. Here, we introduce Regional Expert Networks (REN), the first anatomically-informed MoE framework tailored specifically for medical image classification. REN leverages anatomical priors to train seven specialized experts, each dedicated to distinct lung lobes and bilateral lung combinations, enabling precise modeling of region-specific pathological variations. Multi-modal gating mechanisms dynamically integrate radiomics biomarkers and deep learning (DL) features (CNN, ViT, Mamba) to weight expert contributions optimally. Applied to interstitial lung disease (ILD) classification, REN achieves consistently superior performance: the radiomics-guided ensemble reached an average AUC of 0.8646 $\pm$ 0.0467, a +12.5% improvement over the SwinUNETR baseline (AUC 0.7685, $p=0.031$). Region-specific experts further revealed that lower-lobe models achieved AUCs of 0.88-0.90, surpassing DL counterparts (CNN: 0.76-0.79) and aligning with known disease progression patterns. Through rigorous patient-level cross-validation, REN demonstrates strong generalizability and clinical interpretability, presenting a scalable, anatomically-guided approach readily extensible to other structured medical imaging applications.
@@ I Introduction
In machine learning, Mixture-of-experts (MoE) architectures include multiple expert networks that focus on different aspects of complex data distributions. The main principle underlying MoE success lies in conditional computation; dynamically routing inputs to the most relevant experts while maintaining computational efficiency. However, existing MoE frameworks primarily operate under unconstrained optimization structures that assume homogeneous expert capabilities and domain-agnostic routing mechanisms. We argue that this approach is fundamentally misaligned with medical imaging. Anatomical structure, physiological constraints, and regional disease heterogeneity impose domain-specific specialization requirements that generic routing cannot capture effectively.
Unlike natural image classification, where semantic regions may be arbitrarily distributed, medical images exhibit well-defined anatomical structure where pathological patterns follow predictable regional distributions governed by underlying physiological processes  [1]. Traditional MoE systems cannot leverage these domain-specific priors, leading to suboptimal expert utilization and reduced interpretability, a critical requirement for clinical adoption  [2].
For a specific example in automatic diagnosis of interstitial lung disease (ILD), one can easily exemplify these challenges. ILD encompasses over 200 diverse pulmonary disorders. Accurate diagnosis from high-resolution computed tomography (HRCT) scans is essential, since HRCT plays a central role in distinguishing ILD subtypes and guiding diagnosis in clinical practice [3]. Since each anatomical region may exhibit varying disease manifestations, severity levels, and progression patterns within the same patient, conventional global analysis approaches cannot effectively diagnose with interpretable computation. Current DL methods for ILD classification treat the entire lung as a homogeneous unit, potentially diluting region-specific pathological signals and providing limited interpretability for clinical decision-making [4].
@@ II-B Radiomics and Multi-Modal Integration
Radiomics remains a powerful and actively utilized methodology in medical imaging, particularly when integrated with DL techniques. By extracting quantitative features describing texture, shape, and intensity patterns, radiomics provides complementary information to the hierarchical representations learned by deep networks [13]. This synergy continues to demonstrate significant effectiveness in improving diagnostic accuracy through multi-modal fusion. Traditional machine learning classifiers like XGBoost [14] maintain relevance for high-dimensional radiomics tasks due to their interpretability and capacity to model complex feature interactions. Nevertheless, current fusion strategies often overlook regional anatomical heterogeneity and differential diagnostic importance across tissue regions. Anatomically-informed integration of radiomics with DL represents a promising yet underexplored frontier for enhancing both accuracy and clinical interpretability.
@@ II-C Anatomical Specialization and Regional Analysis
Integrating anatomical priors into DL models enhances performance and interpretability by emulating clinical radiology’s regional examination practices [15]. Regional analysis proves particularly effective for pathologies with anatomically specialized patterns. Conventional approaches that rely on whole-organ analysis often assume anatomical homogeneity and risk overlooking localized variation in pathology [16]. In pulmonary imaging, for example, global feature extraction across entire lung volumes can obscure localized disease signals, reducing both sensitivity and interpretability [17]. This underscores the clinical need for anatomically informed models that capture region-specific patterns and align outputs with radiological workflows [18].
@@ III-A Dataset
Our study utilized a retrospective dataset from the Northwestern Scleroderma Registry comprising 597 patients with 1,898 longitudinal chest CT scans acquired between 2001 and 2023. This study was approved by the Northwestern University Institutional Review Board. All participants provided informed consent for inclusion in the Northwestern Scleroderma Registry at the time of enrollment. The cohort included 489 (81.9%) female patients with a mean age of 63.7 ± 12.7 years (range: 22.1-98.3 years). The study population was predominantly White (479 patients, 80.2%), with 84 (14.1%) Black, 20 (3.4%) Asian, and 14 (2.3%) other race patients. The dataset focused on systemic sclerosis-related conditions with the following disease subtype distribution: limited cutaneous systemic sclerosis (lcSSc): 284 (47.6%), diffuse cutaneous systemic sclerosis (dcSSc): 245 (41.0%), systemic sclerosis sine scleroderma (SSS): 28 (4.7%), with other subtypes (6.7%). Of the total cohort, 365 patients (61.1%) had confirmed ILD, forming the positive class for our binary classification task. Patients contributed an average of 3.2 scans each, reflecting longitudinal monitoring patterns typical in systemic sclerosis care.
@@ III-B Data Preprocessing and Evaluation Strategy
All CT scans, stored as NIFTI files, were processed using identical preprocessing pipelines, accompanied by corresponding lobe segmentation masks and binary classification labels. Lung and lobe segmentation were performed using the state-of-the-art lungmask algorithm [23], which provides robust automated segmentation of pulmonary lobes across diverse imaging protocols and pathological conditions. Our anatomical specialization approach focuses on seven distinct lung regions (automatically determined with lungmask segmentation: the five individual lobes (Left Upper Lobe - LUL, Left Lower Lobe - LLL, Right Upper Lobe - RUL, Right Middle Lobe - RML, Right Lower Lobe - RLL) plus two combined regions (Left Lung combining lobes 1 and 2, Right Lung combining lobes 3, 4, and 5) (Fig. 2). For evaluation, we implement strict patient-level cross-validation to prevent data leakage, where multiple scans per patient could artificially inflate performance if distributed across training and evaluation sets.
@@ III-C Regional Expert Networks (RENs) Architecture
Proposed REN modules specialize in distinct anatomical regions, each processing masked CT inputs corresponding to its assigned lung lobe or lung combination. This design ensures that learning is focused on region-specific pathological patterns rather than global averages. We implement several backbone variants: CNN RENs adopt a 3D convolutional architecture with three progressive convolutional blocks (Conv3D-1, Conv3D-2, Conv3D-3) followed by two fully connected layers (FC-1, FC-2) for binary classification. ViT RENs use a 3D Vision Transformer with patch embedding, positional encoding, multi-head transformer blocks, and a classification head. Mamba RENs apply a state-space model architecture with 3D patch embedding, stacked Mamba blocks with configurable state dimensions, global pooling, and linear classification layers. This multi-architecture design allows comparative evaluation of regional specialization across convolutional, attention-based, and state-space modeling structures.
@@ III-E5 Gating Network Training Configuration
Stage 4 - End-to-End MoE Architecture Training: The final stage trains a unified DL (specifically, we chose SwinUNETR as backbone) mixture-of-experts model for full lung classification, initialized with optimal gating strategies from Stage 3. Rather than training separate regional experts, Stage 4 uses a single end-to-end architecture that processes the entire lung while using anatomical expertise to weight regional contributions.
@@ III-E8 Multi-Component Loss Function
To ensure that the end-to-end model learns not only accurate predictions but also balanced and interpretable expert contributions, we design a multi-component loss function that combines standard classification loss with additional regularizers for gating behavior, weight distribution, and expert diversity. The complete loss includes four components in (28)–(31),
@@
$\mathcal{L}_{\text{gating}}$ $=-\sum_{k=1}^{K}w_{k}\log(w_{k}+\epsilon),$ (29)
@@
III-F Summary of Methodological Innovations
@@ IV-A Stage 2: Individual Expert Performance
IV-B Stage 3: Gating Strategy Evaluation
@@ V Discussion and Concluding Remarks
In conclusion, REN establishes a novel paradigm for anatomically-informed MoE architectures in medical imaging, demonstrating superior ILD classification through lobe-specific expert specialization and multi-modal gating that effectively integrates radiomics and deep learning features. By achieving AUCs of 0.88–0.90 in lower lobes—outperforming clinical benchmarks and traditional DL approaches—REN addresses critical gaps in regional disease heterogeneity, providing interpretable insights via gating weights that align with clinical patterns of basal fibrosis. This framework’s scalability, validated through rigorous patient-level cross-validation, positions it as a transferable solution for other anatomically complex tasks, such as multi-subtype ILD diagnosis or multi-organ segmentation.
@@
We contribute three core methodological innovations. First, cross-architecture analysis of CNN, ViT, and Mamba experts reveals region-specific strengths, suggesting that heterogeneous expert assignment may outperform uniform deployment in anatomically diverse tasks. Second, our evaluation of 11 gating strategies establishes learned gating as consistently superior (1.0–6.2% AUC gain), particularly for ViT, where attention-based mechanisms mitigate domain misalignment. Third, integration of radiomics—handcrafted features rooted in clinical imaging biomarkers—yields a 12.5% performance boost, especially in lower lobes where ILD first manifests. This synergy highlights how radiomics captures subtle, structured signals often missed by end-to-end networks.
@@
Limitations and Future Directions: Our study, while demonstrating strong performance, has several limitations. First, the dataset originates from a single institution and focuses exclusively on systemic sclerosis-related ILD, which may limit generalizability to other ILD subtypes and populations. The 2001–2023 temporal range also introduces variability in imaging protocols and scanner hardware. Additionally, reliance on automated lungmask segmentation may propagate boundary errors throughout the pipeline. From a methodological perspective, our current binary ILD classification does not capture the heterogeneity across ILD subtypes. Expert assignment is anatomically fixed, potentially limiting adaptability to disease-specific spatial patterns. The ensemble’s computational cost—due to training and deploying seven regional networks—may hinder scalability in clinical settings.

### v2

@@ Abstract
Mixture-of-Experts (MoE) architectures have significantly contributed to scalable machine learning by enabling specialized subnetworks to tackle complex tasks efficiently. However, traditional MoE systems lack domain-specific constraints essential for medical imaging, where anatomical structure and regional disease heterogeneity strongly influence pathological patterns. Here, we introduce Regional Expert Networks (REN), the first anatomically-informed MoE framework tailored specifically for medical image classification. REN leverages anatomical priors to train seven specialized experts, each dedicated to distinct lung lobes and bilateral lung combinations, enabling precise modeling of region-specific pathological variations. Multi-modal gating mechanisms dynamically integrate radiomics biomarkers and deep learning (DL) features (CNN, ViT, Mamba) to weight expert contributions optimally. Applied to interstitial lung disease (ILD) classification, REN achieves consistently superior performance: the radiomics-guided ensemble reached an average AUC of 0.8646 $\pm$ 0.0467, a +12.5% improvement over the SwinUNETR baseline (AUC 0.7685, $p=0.031$). Region-specific experts further revealed that lower-lobe models achieved AUCs of 0.88-0.90, surpassing DL counterparts (CNN: 0.76-0.79) and aligning with known disease progression patterns. Through rigorous patient-level cross-validation, REN demonstrates strong generalizability and clinical interpretability, presenting a scalable, anatomically-guided approach readily extensible to other structured medical imaging applications. Code is available on our GitHub https://github.com/NUBagciLab/MoE-REN.
@@ I Introduction
In machine learning, Mixture-of-experts (MoE) architectures include multiple expert networks that focus on different aspects of complex data distributions. The success of MoE architectures stems from conditional computation; dynamically routing inputs to the most relevant experts while maintaining computational efficiency. However, existing MoE frameworks primarily operate under unconstrained optimization structures that assume homogeneous expert capabilities and domain-agnostic routing mechanisms. We argue that this approach is fundamentally misaligned with medical imaging. Anatomical structure, physiological constraints, and regional disease heterogeneity impose domain-specific specialization requirements that generic routing cannot capture effectively.
Unlike natural image classification, where semantic regions may be arbitrarily distributed, medical images exhibit well-defined anatomical structure where pathological patterns follow predictable regional distributions governed by underlying physiological processes [3, 14]. Traditional MoE systems cannot leverage these domain-specific priors, leading to suboptimal expert utilization and reduced interpretability, a critical requirement for clinical adoption [30].
For a specific example in automatic diagnosis of interstitial lung disease (ILD), one can easily exemplify these challenges. ILD encompasses over 200 diverse pulmonary disorders. Accurate interpretation of high-resolution computed tomography (HRCT) scans is essential for distinguishing ILD subtypes and guiding clinical management [22]. Since each anatomical region may exhibit varying disease manifestations, severity levels, and progression patterns within the same patient, conventional global analysis approaches struggle to provide both accurate diagnosis and interpretable outputs. Current DL methods for ILD classification treat the entire lung as a homogeneous unit, potentially diluting region-specific pathological signals and providing limited interpretability for clinical decision-making [13].
@@ II-B Radiomics and Multi-Modal Integration
Radiomics remains a powerful methodology in medical imaging, particularly when integrated with DL techniques. By extracting quantitative features describing texture, shape, and intensity patterns, radiomics provides complementary information to the hierarchical representations learned by deep networks [10]. This synergy continues to demonstrate significant effectiveness in improving diagnostic accuracy through multi-modal fusion. Traditional machine learning classifiers like XGBoost [4] maintain relevance for high-dimensional radiomics tasks due to their interpretability and capacity to model complex feature interactions. Nevertheless, current fusion strategies often overlook regional anatomical heterogeneity and differential diagnostic importance across tissue regions [32]. Anatomically-informed integration of radiomics with DL represents a promising yet underexplored frontier for enhancing both accuracy and clinical interpretability.
@@ II-C Anatomical Specialization and Regional Analysis
Integrating anatomical priors into DL models enhances performance and interpretability by emulating clinical radiology’s regional examination practices [7, 17]. Regional analysis proves particularly effective for pathologies with anatomically specialized patterns [2]. Conventional approaches that rely on whole-organ analysis often assume anatomical homogeneity and risk overlooking localized variation in pathology [20, 33]. In pulmonary imaging, for example, global feature extraction across entire lung volumes can obscure localized disease signals, reducing both sensitivity and interpretability [26, 9]. This underscores the clinical need for anatomically informed models that capture region-specific patterns and align outputs with radiological workflows [21].
@@ III-A Dataset
Our study utilized a retrospective dataset from the Northwestern Scleroderma Registry comprising 597 patients with 1,898 longitudinal chest CT scans acquired between 2001 and 2023. This study was approved by the Northwestern University Institutional Review Board. All participants provided informed consent for inclusion in the Northwestern Scleroderma Registry at the time of enrollment. The cohort included 489 (81.9%) female patients with a mean age of 63.7 ± 12.7 years (range: 22.1-98.3 years). The study population was predominantly White (479 patients, 80.2%), with 84 (14.1%) Black, 20 (3.4%) Asian, and 14 (2.3%) other race patients. The dataset focused on systemic sclerosis-related conditions with the following disease subtype distribution: limited cutaneous systemic sclerosis (lcSSc): 284 (47.6%), diffuse cutaneous systemic sclerosis (dcSSc): 245 (41.0%), systemic sclerosis sine scleroderma (SSS): 28 (4.7%), with other subtypes (6.7%). Of the total cohort, 365 patients (61.1%) had confirmed ILD (via multidisciplinary consensus), forming the positive class for our binary classification task. Patients contributed an average of 3.2 scans each, reflecting longitudinal monitoring patterns typical in systemic sclerosis care.
@@ III-B Data Preprocessing and Evaluation Strategy
All CT scans, stored as NIFTI files, were processed using identical preprocessing pipelines, accompanied by corresponding lobe segmentation masks and binary classification labels. Lung and lobe segmentation were performed using the state-of-the-art lungmask algorithm [12], which provides robust automated segmentation of pulmonary lobes across diverse imaging protocols and pathological conditions. Our anatomical specialization approach focuses on seven distinct lung regions (automatically determined with lungmask segmentation): the five individual lobes (Left Upper Lobe - LUL, Left Lower Lobe - LLL, Right Upper Lobe - RUL, Right Middle Lobe - RML, Right Lower Lobe - RLL) plus two combined regions (Left Lung combining lobes 1 and 2, Right Lung combining lobes 3, 4, and 5) (Fig. 2). For evaluation, we implement strict patient-level cross-validation to prevent data leakage, where multiple scans per patient could artificially inflate performance if distributed across training and evaluation sets.
@@ III-C Regional Expert Networks (RENs) Architecture
Proposed REN modules specialize in distinct anatomical regions, each processing masked CT inputs corresponding to its assigned lung lobe or lung combination. This design ensures that learning is focused on region-specific pathological patterns rather than global averages. We implement several backbone variants: CNN RENs adopt a 3D convolutional architecture with three progressive blocks (24→48→96 filters, $3\times 3\times 3$ kernels), each followed by batch normalization, ReLU activation, and $2\times 2\times 2$ max pooling, before two fully connected layers (256→64 units) with dropout 0.5. ViT RENs partition volumes into $8\times 8\times 8$ patches, embed them into 384 dimensions, and process through 6 transformer blocks (8 attention heads, MLP ratio 4.0, GELU activation) with learnable positional encodings. Mamba RENs similarly use $8\times 8\times 8$ patches (256-dimensional embeddings) processed through 4 Mamba blocks ($d_{\text{state}}=64$, $d_{\text{conv}}=4$, expansion factor 2) with SiLU activation. This multi-architecture design allows comparative evaluation of regional specialization across convolutional, attention-based, and state-space modeling structures.
@@ III-E Training Pipeline
Stage 3 – Staged Ensemble with Fixed Gating. In Stage 3, regional experts are combined using gating weights derived from validation data, including validation-AUC–based weighting and learned gating networks. Crucially, once selected, these gating weights are fixed and are not updated during subsequent inference or training. This staged design decouples expert learning from expert combination and preserves region-specific specialization.
@@
For each fold and architecture, gating strategies were evaluated exclusively on the validation split. The optimal gating function was selected as the one maximizing validation AUC and then fixed for testing, ensuring no test-set leakage.
Definition of Ensemble Variants: For clarity and reproducibility, we explicitly define all ensemble configurations used throughout this study:
  • Radiomics-Guided: Expert weights are derived exclusively from validation AUCs of radiomics (XGBoost) regional models and applied to deep expert outputs.
  • Validation-Weighted: Expert outputs are combined using fixed, global weights derived from normalized validation AUCs of the corresponding deep models.
  • CNN / ViT / Mamba Gated MoE: End-to-end MoE models in which expert weights are input-dependent and learned via gating networks that are jointly optimized with the backbone.
All reported results correspond to Tables I and II.
@@ III-E1 Enhanced AUC Normalization
In sparsemax normalization, the threshold $\tau$ is computed as the unique value that ensures $\sum_{k}\max(\text{AUC}_{k}^{\text{val}}-\tau,0)=1$, following the standard sparsemax formulation.
@@ III-E5 Gating Network Training Configuration
Stage 4 - End-to-End MoE Architecture Training: Stage 4 initializes expert and gating parameters using the optimal weights identified in Stage 3 and then jointly optimizes the backbone, experts, and gating in an end-to-end manner. Unlike Stage 3, which preserves independent regional training, Stage 4 employs a unified architecture that processes the full lung volume while adaptively weighting anatomically constrained expert contributions.
SwinUNETR was selected as the Stage 4 backbone due to its strong and well-documented performance in volumetric medical imaging, its ability to balance global contextual modeling with local spatial fidelity [1], and its widespread clinical adoption. The hierarchical Swin Transformer encoder captures long-range dependencies across lung regions, while the U-Net–style decoder preserves spatial structure, making it well suited for aggregating region-specific expert evidence into a unified representation. In our baseline experiments, SwinUNETR achieved higher full-lung AUC than alternative CNN-, ViT-, and Mamba-based backbones, motivating its use as the most competitive single-model reference. In addition, SwinUNETR benefits from large-scale medical imaging pretraining, which improves optimization stability and generalization in limited-data regimes. Importantly, REN is backbone-agnostic, and SwinUNETR is used here to provide a robust and reproducible reference for evaluating the proposed anatomically informed MoE formulation rather than to introduce architectural novelty.
@@ III-E8 Multi-Component Loss Function
To ensure that the end-to-end model learns not only accurate predictions but also balanced and interpretable expert contributions, we design a multi-component loss function that combines standard classification loss with additional regularizers for entropy behavior, weight distribution, and expert diversity. The complete loss includes four components in (28)–(31),
@@
$\mathcal{L}_{\text{entropy}}$ $=-\sum_{k=1}^{K}w_{k}\log(w_{k}+\epsilon),$ (29)
@@
III-F Scope of Anatomical Priors
We operationalize anatomical knowledge as an inductive bias through region-constrained expert assignment, rather than as a formal probabilistic prior. Specifically, expert specialization is enforced through anatomically defined region decomposition and masked inputs, reflecting radiological practice.
III-G Summary of Methodological Innovations
@@
Role of Radiomics in REN: We acknowledge that radiomics–deep learning fusion has been explored in prior ILD and thoracic imaging studies [22]. In our current study, our contribution lies in how radiomics is used within REN: radiomics features are not fused at the representation level, but serve as region-specific reliability signals that guide expert weighting within an anatomically constrained MoE framework. This design preserves modularity between handcrafted and learned representations while enabling region-aware interpretability through anatomically grounded expert weighting.
@@ IV-A Stage 2: Individual Expert Performance
IV-B Stage 3: Learned and Gating Strategy Evaluation
@@ IV-C Stage 4: End-to-End Integration Results
TABLE III: Training Complexity and Resource Requirements Across Models for Each Fold
Model Training Time Peak Memory Parameters
(GPU-hours) (GB) (Millions)
CNN Expert (single region) 0.82 1.94 42.6
ViT Expert (single region) 0.82 1.66 0.6
Mamba Expert (single region) 0.81 0.27 0.7
SwinUNETR Baseline 0.91 12.98 62.2
MoE Pipeline (CNN expert) 5.10 19.39 63.6
MoE Pipeline (ViT expert) 4.98 19.39 63.6
MoE Pipeline (Mamba expert) 5.04 19.39 63.6
MoE Pipeline (Radiomics expert) 5.07 19.39 63.6
Gated MoE (end-to-end) 5.02 19.74 64.7
TABLE IV: Ablation Study of Gating Regularization Effects on Expert Behavior of all Experts
Configuration Mean Active Usage Mean Cosine
Experts Variance Entropy Similarity
Baseline (Full Loss) 3.10 237.2 1.43 0.49
No Regularization 2.25 920.2 0.85 0.83
Remove Entropy 2.58 620.2 1.00 0.41
Remove Load Balancing 3.01 290.3 1.35 0.53
Remove Diversity 3.34 183.2 1.49 0.82
Entropy Weight = 0.001 2.80 532.8 1.12 0.44
Entropy Weight = 0.002 2.69 673.4 1.07 0.52
Entropy Weight = 0.01 3.76 97.5 1.68 0.54
Entropy Weight = 0.02 5.12 60.7 1.79 0.60
Load Balance Weight = 0.001 3.36 238.0 1.44 0.52
Load Balance Weight = 0.005 3.49 142.2 1.58 0.48
Load Balance Weight = 0.02 3.48 226.1 1.45 0.57
Load Balance Weight = 0.05 3.11 305.4 1.38 0.61
Diversity Weight = 0.001 3.03 223.6 1.38 0.83
Diversity Weight = 0.005 3.05 256.7 1.40 0.60
Diversity Weight = 0.02 2.61 572.2 1.21 0.38
Diversity Weight = 0.05 2.52 721.3 1.12 0.29
@@
Detailed training time, memory usage, and parameter counts for all expert and MoE variants are reported in Table III. Although staged MoE pipelines incur higher total training cost due to independent expert training, peak memory usage remains below 20 GB and per-model training time is comparable to the SwinUNETR baseline, supporting practical clinical deployment.
@@
Ablation of Gating Regularization Terms: Table IV analyzes the effect of individual gating regularizers on expert utilization and redundancy. Removing all regularization leads to severe expert collapse, characterized by high usage variance and strong feature similarity between experts. Entropy regularization controls the number of active experts and prevents overconfident routing, while load-balancing regularization reduces expert usage variance and promotes equitable expert participation. Diversity regularization plays a critical role in reducing redundancy, as removing it substantially increases cosine similarity between expert representations. Together, these components stabilize expert behavior, balancing specialization, coverage, and interpretability.
@@ V Discussion and Concluding Remarks
In conclusion, REN demonstrates a promising approach for anatomically-informed MoE architectures in medical imaging, demonstrating superior ILD classification through lobe-specific expert specialization and multi-modal gating that effectively integrates radiomics and deep learning features. By achieving AUCs of 0.88–0.90 in lower lobes—outperforming clinical benchmarks and traditional DL approaches—REN addresses critical gaps in regional disease heterogeneity, providing interpretable insights via gating weights that align with clinical patterns of basal fibrosis. This framework’s scalability, validated through rigorous patient-level cross-validation, positions it as a transferable solution for other anatomically complex tasks, such as multi-subtype ILD diagnosis or multi-organ segmentation.
@@
Our results reveal three key insights. First, cross-architecture analysis of CNN, ViT, and Mamba experts reveals region-specific strengths, suggesting that heterogeneous expert assignment may outperform uniform deployment in anatomically diverse tasks. Second, our evaluation of 11 gating strategies establishes learned gating as consistently superior (1.0–6.2% AUC gain), particularly for ViT, where attention-based mechanisms mitigate domain misalignment. Third, integration of radiomics—handcrafted features rooted in clinical imaging biomarkers—yields a 12.5% performance boost, especially in lower lobes where ILD first manifests. This synergy highlights how radiomics captures subtle, structured signals often missed by end-to-end networks.
@@
Inductive Bias and Data-Regime Effects in ViT and Mamba Experts: The underperformance of ViT and Mamba experts in our experiments is unlikely to reflect architectural inferiority, but rather a mismatch between their inductive biases and the available data regime. Both architectures are optimized for large-scale training and benefit from modeling long-range global context, which may be less critical for ILD diagnosis where discriminative signals are often localized fibrotic textures and density changes in 3D CT. In contrast, CNN-based models and radiomics features explicitly emphasize local spatial patterns and intensity statistics, making them more effective under limited sample sizes. This interpretation is consistent with prior findings in medical imaging, where transformer-style architectures typically require substantially larger datasets or extensive pretraining to consistently outperform convolutional baselines.
Synergistic Role of Radiomics and Anatomical Specialization: The 12.5% performance gain observed with radiomics-guided weighting does not arise from radiomics features alone. Radiomics models trained on whole-lung regions perform substantially worse than lobe-specific radiomics experts, indicating that anatomical specialization is a necessary condition for effective radiomics utilization. Conversely, anatomically specialized deep models without radiomics guidance do not achieve comparable gains. Together, these results indicate that the improvement arises from the synergy between anatomically localized feature extraction and pathology-aware radiomics descriptors, rather than from either component in isolation.
@@
Clinical Deployment Considerations. REN is intended as a decision-support system that augments radiologist interpretation by providing anatomically grounded, region-aware evidence rather than fully automated diagnosis.
Dense Expert Activation and Clinical Trade-offs. Unlike traditional MoE architectures that emphasize sparse expert activation for computational efficiency, REN intentionally activates all regional experts. This design choice reflects clinical radiological practice, where radiologists systematically assess all anatomically relevant lung regions to ensure comprehensive disease evaluation. Rather than enforcing hard sparsity, REN introduces soft sparsity through gating weights that downweight less informative regions on a per-patient basis. This approach trades strict computational sparsity for anatomical completeness, interpretability, and clinical trustworthiness, while maintaining computational feasibility for clinical deployment.
Model Compression and Deployment: REN is well-suited to compression due to its staged design. A practical deployment strategy is knowledge distillation, where the trained REN ensemble serves as a teacher and a single student network is optimized using both soft prediction targets and region-level intermediate features. Additional strategies include expert pruning based on consistently low gating weights, top-$k$ regional routing to reduce inference cost, and parameter sharing across experts with region-specific heads. Standard structured pruning and post-training quantization can further reduce memory and latency without degrading diagnostic performance.
To facilitate clinical integration, the proposed anatomically informed MoE model can be deployed as a decision-support module that processes routine chest CT scans, performs automated lung and regional analysis, and outputs a global ILD severity label with calibrated confidence and regional expert contributions. The MoE framework inherently supports interpretability by indicating which anatomical experts influence the final decision, enabling clinically aligned review (e.g., per-lobe summaries). Practical deployment requires robust handling of protocol variability and common confounders such as infection, edema, motion, and emphysema, along with quality-control checks and a human-in-the-loop workflow. Before clinical use, multi-center external validation and prospective evaluation are recommended, followed by monitoring for dataset shift, version control, and governance aligned with privacy and regulatory standards. Accordingly, REN provides calibrated, region-level evidence that supports clinical interpretation and review, rather than replacing radiologist judgment.
@@
Limitations and Future Directions: Our study, while demonstrating strong performance, has several limitations. First, the dataset originates from a single institution and focuses exclusively on systemic sclerosis-related ILD, which may limit generalizability to other ILD subtypes and populations. The 2001–2023 temporal range also introduces variability in imaging protocols and scanner hardware. Additionally, reliance on automated lungmask segmentation may propagate boundary errors throughout the pipeline. Future work will evaluate sensitivity to segmentation quality through controlled perturbation experiments. From a methodological perspective, our current binary ILD classification does not capture the heterogeneity across ILD subtypes. Expert assignment is anatomically fixed, potentially limiting adaptability to disease-specific spatial patterns. The ensemble’s computational cost—due to training and deploying seven regional networks—may hinder scalability in clinical settings.
The cohort originates from a single academic health system, however it includes patients imaged across 11 affiliated hospitals using a wide range of scanner vendors, models, and acquisition protocols. This heterogeneity reduces the likelihood of systematic bias arising from a single imaging device or protocol. While referral patterns at a tertiary care center may still enrich for more complex cases, the diversity of imaging sources within the Northwestern network mitigates substantial institutional bias in learned feature distributions. Ongoing work includes planned multi-institutional validation through collaborations with external academic centers. These efforts aim to evaluate REN across heterogeneous scanners, patient demographics, and ILD etiologies, and to assess the robustness of regional gating behavior under domain shift. While cross-institutional data sharing poses inherent logistical and regulatory challenges, our team is actively working toward extending this study into a multi-center validation framework.
Although this study emphasizes ILD severity classification, the anatomically-informed MoE design is not necessarily specific to subtypes. Since each expert functions within anatomically constrained regions, the framework should remain applicable across ILD subtypes sharing common spatial and structural factors that influence physiology (e.g., extent of fibrosis, traction bronchiectasis, honeycombing, ground-glass opacities) but have different distribution patterns. However, performance may vary by subtype due to differences in typical regional involvement (e.g., basal/subpleural dominance in UIP/IPF, mid–upper lung involvement in sarcoidosis or hypersensitivity pneumonitis, peribronchovascular patterns in connective tissue disease–ILD), variations in prevalence and label noise, and confounding coexisting processes (infection, edema, emphysema). Therefore, applying the model to other subtypes should be validated with external testing stratified by subtype, and—where labels allow—either (i) recalibration, (ii) expert fine-tuning, or (iii) multi-task training that jointly predicts subtype and severity.